1/1

Mèng Hàorán,  Mèng Hàorán,
The glow on western mountains  Shān guāng

tradukita de N. N. 81

Xià rì nán tíng huái xīn dà

The glow on western mountains quickly sets, The moon is climbing over the eastern lake. My hair loose, I enjoy the evening cool, I lie in peace before the open window. The wind spreads lotus scent all through the air, The sound of dripping bamboo dew is clear. Although I’d like to fetch my qin and play, To my regret, there is no-one to hear. So touched by this, I think of my old friend, Throughout the night, I’m troubled by my dreams.

Shān guāng hū xī luò Chí yuè jiàn dōng shàng Sǎn fà chéng xī liáng Kāi xuān wò xián chǎng Hé fēng sòng xiāng qì Zhú lù dī qīng xiǎng Yù qǔ míng qín tán Hèn wú zhī yīn shǎng Gǎn cǐ huái gù rén Zhōng xiāo láo mèng xiǎng

Traduko de la ĉina poemo “Shān guāng” de Mèng Hàorán (∗689 – †740) en la Anglan de N. N. 81.

Arg-995-2010 (2014-04-20 00:35:06)

Tiu ĉi laŭsenca angligo troviĝas en http: // www. chinese-poems. com/ m6. html .

Verkinto de tiu ĉi ĉina poemo estas Mèng Hàorán (∗689 – †740).

Arg-995-2006 (2014-04-21 20:01:57)

La ĉina-lingva teksto de tiu ĉi poemo troveblas sub http: // www. chinese-poems. com/ m6. html . Tie estas indikita ankaŭ laŭvorta kaj laŭsenca angligoj. Pri la aŭtoro vidu la vikipediejon http: // en. wikipedia. org/ wiki/ Meng_Haoran .

http://www.poezio.net/version?poem-id=995&version-id=,2010,2006